A store has 84 blouses of three sizes. If 1/3 are small sized,1/4 are small sized, and the rest are large sized, how many blouses are large sized?
Found 3 solutions by dkppathak, ikleyn, josgarithmetic:
Answer by dkppathak(439)

A store has 84 blouses of three sizes. If 1/3 are small sized,1/4 are small sized, and the rest are large sized, how many blouses are large sized?
0 solutions
small +very small size=1/3+1/4=7/12
large size=1-7/12=5/12
large size =84x5/12=35
large size 35

Answer by ikleyn(52781)

.

The problem is posed in an absurdist way.

It does not deserve any consideration.

It only deserves to be thrown to the  TRACH  BIN.

The author deserves to be ticketed.



Answer by josgarithmetic(39617)

I suspect the problem should have said, "If 1/3 are small sized,1/4 are medium sized, and the rest are large sized,"